During a time interval of 90 minutes a car


(a) Runs at 50 km/hr for the first 20 min,


(b) Accelerates uniformly to 90 km/hr for the next 10 min,


(e) Runs at the speed of 90 km/hr in the next 40 min, and (d) Decelerates uniformly to a stop in the remaining 20 min.


Calculate the (a) acceleration (b) deceleration (e) average speed during entire period.


1
Expert's answer
2022-02-20T08:20:20-0500

a)

"a_1=\frac{v_2-v_1}{t_2}=0.0185~\frac m{c^2},"

b)

"a_2=\frac{v_2}{t_4}=0.0208~\frac m{c^2},"

c)

"‹v›=\frac{s_1+s_2+s_3+s_4}{t_1+t_2+t_3+t_4}=68.9~\frac{km}h."